OpenStudy (anonymous):

how do you solve 9^6a <31

OpenStudy (anonymous):

\[\LARGE 9^{6a}<31\] \[\LARGE \log(9)^{6a}<\log(31)\] \[\LARGE 6a\cdot \log(9)<\log(31)\] \[\LARGE 6a<\frac{\log(31)}{\log(9)}\] \[\LARGE 6a<\log_9(31 )\] \[\LARGE a<\frac{\log_9(31)}{6}\] UGLY ANSWER ! :( isn't it?